# Dissecting Lepton Number Violation in the Left-Right Symmetric Model: $0\nu\beta\beta$ decay, Møller scattering, and collider searches

### v1

@@ VI Conclusion
We employ the effective field theory framework to calculate the ${0\nu\beta\beta}$-decay rate with the contributions arising from the mediation of right-handed neutrinos and the right-handed doubly-charged scalar, considering possible non-zero $W_{L}-W_{R}$ mixing. Compared to the earlier work of Ref. Dev et al. (2018), the effective field theory framework provides a more realistic assessment of the ${0\nu\beta\beta}$-decay sensitivity and yields generally significantly greater parameter space reach.
@@
Acknowledgements.
GL is supported by the National Natural Science Foundation of China under Grant No. 12347105, the Guangdong Basic and Applied Basic Research Foundation (2024A1515012668), Fundamental Research Funds for the Central Universities in Sun Yat-sen University (23qnpy62), and SYSU startup funding. MJRM and SUQ were supported in part under the US Department of Energy contract DE-SC0011095. SUQ also acknowledges the support by the U.S. DOE under Grant No. DE-FG02-00ER41132. SUQ would like to thank the High Energy Theory Group at William & Mary for their hospitality, where part of this work was conducted.

### v2

@@ 6 Conclusion
We employ the effective field theory framework to calculate the $0\nu\beta\beta$}-decay rate with the contributions arising from the mediation of right-handed neutrinos and the right-handed doubly-charged scalar, considering possible non-zero $W_{L}-W_{R}$ mixing. Compared to the earlier work of Ref. [Dev:2018sel], the effective field theory framework provides a more realistic assessment of the $0\nu\beta\beta$}-decay sensitivity and yields generally significantly greater parameter space reach.
@@
Acknowledgments
GL is supported by the National Natural Science Foundation of China under Grant No. 12347105, the Guangdong Basic and Applied Basic Research Foundation (2024A1515012668), and SYSU startup funding. MJRM and SUQ were supported in part under the US Department of Energy contract DE-SC0011095. SUQ also acknowledges the support of the U.S. DOE under Grant No. DE-FG02-00ER41132. SUQ would like to thank the High Energy Theory Group at William & Mary for their hospitality, where part of this work was conducted.